OpenStudy (anonymous):

when the temperature is 68°F, the speed of sound in air is approximately 343 meters per second. Approximately how many kilometers does sound travel in five seconds?

OpenStudy (superdavesuper):

distance in 5 seconds = distance in 1 second x 5 right?

OpenStudy (anonymous):

idk that's why I am asking

OpenStudy (anonymous):

@johnweldon1993

OpenStudy (superdavesuper):

hmm the eqn in my last post is the right one to solve this prob....

OpenStudy (johnweldon1993):

Well we need a conversion here first... Notice what they give us...343 METERS per seconds.. Then they ask for KILOMETERS... How many meters are in 1 Kilometer?

OpenStudy (anonymous):

0.001

OpenStudy (superdavesuper):

thanks @johnweldon1993 i missed that!

OpenStudy (anonymous):

ops wait hold on

OpenStudy (anonymous):

correction=1000

OpenStudy (johnweldon1993):

Well 1000M = 1 KM yes but you were right the first time...we need the .001 because we want to know how many KM are in 1 M So.. \[\large \frac{343Meters}{Second} \times \frac{.001 KM}{1 Meter} = \frac{.343KM}{Second}\] Now just multiply that by 5

OpenStudy (anonymous):

1716.72 @johnweldon1993

OpenStudy (superdavesuper):

plz read john's eqn again....carefully :)

OpenStudy (johnweldon1993):

@Xx_J.L.S_Xx so we have \[\large .343KM/sec \times 5\]

OpenStudy (anonymous):

1.715

OpenStudy (johnweldon1993):

Looks right to me :) 1.715km/sec

OpenStudy (anonymous):

what next?

OpenStudy (johnweldon1993):

*That's the answer :)

OpenStudy (anonymous):

omg thanks!

OpenStudy (johnweldon1993):

Anytime :)